# ArXiv Preprint

### v1

@@ 4.1 Experimental Setup
Each narrative (approximately 100 words) is decomposed into three time frames via prompting. For each frame, we extract three predicate triples, which are matched to GameTileNet assets using semantic embeddings and placed within procedurally generated terrains.
Table 3 summarizes the evaluation dataset.
@@ 4.2 Tile Matching Accuracy
We first examine whether the tiles selected by semantic matching correspond well to the narrative objects. Evaluation considers (1) whether the top-1 matched tile is semantically appropriate, (2) whether the assigned affordance matches the expected gameplay role, and (3) whether the system produces a diverse set of tiles across each story.
Per-story results are shown in Table 4, and aggregate results across all 30 scenes are given in Table 5.
@@ 4.3 Spatial Predicate Satisfaction
Predicate satisfaction averaged 72% across all stories (Table LABEL:tab:spatial), with the majority of scenes achieving two-thirds or more of their relational constraints. Stories 6 and 1 achieved the highest consistency (89% and 78%), while Story 10 lagged (56%), typically due to conflicting placement constraints or lack of sufficient map space. These results suggest that the procedural layout is capable of enforcing basic spatial relations but can be brittle when multiple constraints interact. Improvements such as constraint-solving or affordance-informed placement could further enhance satisfaction rates.
@@ 5.1 Strengths and Generalization
The evaluation highlights several promising aspects of our approach. First, semantic alignment between narrative descriptions and visual tiles is stable across all ten stories (Table 5), indicating that embedding-based retrieval provides a reliable foundation for mapping open-ended narrative text into game assets. Second, the system maintains high tile diversity (mean 0.92), suggesting that it can produce varied outputs without excessive repetition, an important property for replayability and player engagement. Third, spatial predicate satisfaction averaged 72% (Table LABEL:tab:spatial), demonstrating that even a lightweight rule-based layout generator can enforce a substantial fraction of narrative constraints. Together, these findings suggest that the pipeline generalizes across different story contexts, making it adaptable for varied game scenarios.
@@ 5.2 Limitations
Despite these strengths, several limitations remain. Affordance matching showed high variance (0.27-0.55), pointing to a gap between visual similarity and gameplay semantics. This limitation partly arises from coverage issues in the GameTileNet dataset: some narrative objects (e.g., "lantern," "archway") lack sufficiently representative tiles, forcing the system to select approximations. Moreover, while semantic embeddings capture descriptive similarity, they do not account for functional roles (e.g., terrain vs. collectible), leading to frequent misplacements.
Another limitation is that symbolic layers were not fully leveraged for spatial reasoning. Our current layout engine checks constraints but does not resolve conflicts, which leads to brittle performance when multiple spatial predicates interact. Finally, while narrative-to-scene generation is inherently open-ended, this openness introduces evaluation challenges. Metrics such as diversity and cosine similarity are not only influenced by the correct interpretation of narrative objects, but also by the comprehensiveness of the tile set and the inherent ambiguity of natural language descriptions.
@@ 6 Conclusion
We presented a pipeline for generating game scenes from narrative text by aligning LLM-derived predicates with the GameTileNet dataset and rendering layered maps using procedural terrain generation. Our evaluation across ten stories demonstrated that semantic matching provides stable visual alignment with narrative objects, while affordance alignment and spatial relation enforcement remain challenging. These findings highlight both the promise of semantic embeddings for bridging text and assets and the need for deeper affordance-aware reasoning to ensure gameplay consistency.
This work provides an early step toward narrative-driven procedural content generation. Future directions include integrating symbolic reasoning for more reliable spatial and temporal coordination, expanding the coverage of tile datasets to reduce gaps in representation, and supporting interactive or co-creative workflows where designers and players can iteratively refine generated scenes. We see these developments as important next steps toward practical tools that blend narrative expression with playable game environments.
@@ Appendix A Online Resources
The Code and examples are released publicly via GitHub:

### v2

@@ 4.1 Experimental Setup
Each narrative (approximately 100 words) is decomposed into three time frames via prompting. For each frame, we extract three predicate triples, which are matched to GameTileNet assets using semantic embeddings and placed within procedurally generated terrains. Table 3 summarizes the evaluation dataset.
@@ 4.2 Tile Matching Accuracy
We first examine whether the tiles selected by semantic matching correspond well to the narrative objects. Evaluation considers (1) whether the top-1 matched tile is semantically appropriate, (2) whether the assigned affordance matches the expected gameplay role, and (3) whether the system produces a diverse set of tiles across each story. Per-story results are shown in Table 4, and aggregate results across all 30 scenes are given in Table 5.
@@ 4.3 Spatial Predicate Satisfaction
Predicate satisfaction averaged 72% across all stories (Table 4), with the majority of scenes achieving two-thirds or more of their relational constraints. Stories 6 and 1 achieved the highest consistency (89% and 78%), while Story 10 lagged (56%), typically due to conflicting placement constraints or lack of sufficient map space. These results suggest that the procedural layout is capable of enforcing basic spatial relations but can be brittle when multiple constraints interact. Improvements such as constraint-solving or affordance-informed placement could further enhance satisfaction rates.
@@ 5.1 Strengths and Generalization
The evaluation highlights several promising aspects of our approach. First, semantic alignment between narrative descriptions and visual tiles is stable across all ten stories (Table 5), indicating that embedding-based retrieval provides a reliable foundation for mapping open-ended narrative text into game assets. Second, the system maintains high tile diversity (mean 0.92), suggesting that it can produce varied outputs without excessive repetition, an important property for replayability and player engagement. Third, spatial predicate satisfaction averaged 72% (Table 4), demonstrating that even a lightweight rule-based layout generator can enforce a substantial fraction of narrative constraints. Together, these findings suggest that the pipeline generalizes across different story contexts, making it adaptable for varied game scenarios.
@@ 5.2 Limitations
Despite these strengths, several limitations remain. Affordance matching showed high variance (0.27–0.55), revealing a gap between visual similarity and gameplay semantics. This partly stems from limited coverage in the GameTileNet dataset: some objects (e.g., "lantern," "archway") lack representative tiles, forcing approximate matches. Semantic embeddings also capture descriptive but not functional similarity (e.g., terrain vs. collectible), leading to occasional misplacements. Symbolic layers were not fully leveraged for spatial reasoning, the layout engine checks constraints but does not resolve conflicts, causing brittle performance when multiple predicates interact. Finally, narrative-to-scene generation is inherently open-ended, complicating evaluation. Metrics such as diversity and cosine similarity depend not only on correct object interpretation but also on tile coverage and the ambiguity of natural language descriptions. The current three-frame segmentation follows a fixed narrative template; future versions will explore data-driven segmentation that adapts frame count to story complexity. Ablation studies on individual modules (terrain generation, semantic matching, spatial refinement) were not conducted; future work will quantify their contributions to overall scene coherence.
@@ 6 Conclusion
We presented a pipeline for generating game scenes from narrative text by aligning LLM-derived predicates with the GameTileNet dataset and rendering layered maps using procedural terrain generation. Our evaluation across ten stories demonstrated that semantic matching provides stable visual alignment with narrative objects, while affordance alignment and spatial relation enforcement remain challenging. These findings highlight both the promise of semantic embeddings for bridging text and assets and the need for deeper affordance-aware reasoning to ensure gameplay consistency. This work provides an early step toward narrative-driven procedural content generation. Future directions include integrating symbolic reasoning for more reliable spatial and temporal coordination, expanding the coverage of tile datasets to reduce gaps in representation, and supporting interactive or co-creative workflows where designers and players can iteratively refine generated scenes. We see these developments as important next steps toward practical tools that blend narrative expression with playable game environments.
Declaration on Generative AI
Generative AI tools were used only to assist with language refinement and LaTeX formatting under the authors’ direction. All conceptual contributions, design, implementation, and analysis were produced by the authors.
@@ Appendix A Online Resources
The Code and examples are released publicly via https://github.com/RimiChen/2025_NarrativeScene.